ms.wolk is approaching a stoplight moving with a velocity of 30.0 m/s. the light turns yellow, and wolk applies the breaks and skids to a stop. if wolks deceleration is 8.00 m/s2, determine the displacement of the car during the skidding process.
1
Expert's answer
2012-10-08T11:18:37-0400
Let's find the time elapsed by skidding process:

V = at ==> t = V/a = 30.0[m/s]/8.00[m/s²] = 3.75[s].

Now we can determine the displacement of the car:

D = Vt - at²/2 = (V - at/2)t = (30.0[m/s] - 8.00[m/s²]*3.75[s]/2)*3.75[s] = 56.25[m].

So, the displacement of the car during the skidding process was 56.25[m].